Fruitland is a closed economy that always produces and consumes 50 apples and 50 bananas. Initially both apples and bananas cost $1 each. Then demand shifts, so the price of apples doubles to $2 and the price of bananas halves to 50 cents, but production and consumption of apples and bananas stays the same.

Calculate the effect on: Nominal GDP, Real GDP, the GDP Deflator, and the Consumer Price Index.
Answer #1

Nominal GDP = P*Q

Real GDP = Base price X Q

GDP deflator =( NGDP/ RGDP)*100

Year Quantity of apples Quantity of bananas Price of apples(in $) Price of bananas(in $) NGDP RGDP GDP deflator
Initial 50 50 1 1

50*1+50*1

= $100

50*1+50*1

=$100

(100/100)*100

=100

Next year 50 50 2 0.5

50*2+0.5*50

=125$

50*1+50*1

=100$

(125/100)*100

=125

Cost of basket in initial year = 50*1+50*1 = 100

Cost of basket after price change = 50*2+50*0.5 = 125

CPI = (125/100)*100 = 125